# Fine-Grained Domain Generalization with Feature Structuralization

### v1

@@ VI Conclusion
In this paper, we assimilate insights from cognitive psychology to advance a feature structuralization approach to tackle the fine-grained domain generalization challenge. The disentanglement of learned features into common, specific, and confounding segments, the feature alignment constraints facilitated by three loss functions, and the label alignment technique collectively contribute to the creation of the feature structuralized domain generalization model. Semantic constraints and alignments are then applied through the incorporation of multi-granularity structural knowledge. The experimental results on three benchmarks demonstrate a notable enhancement in FGDG performance and explainability.
Despite FSDG’s systemic efficacy, it brings forth new challenges and prompts intriguing questions. For example, the pre-constructed granularity structure is used to achieve semantic feature alignments. However, some datasets may lack granularity structures, posing challenges to granularity constraints. Notably, one of the prevailing trends in AI research is the pursuit of large-scale datasets, leading to a significant increase in the number of data categories [55]. Such large-scale datasets inherently exhibit hierarchical multi-granularity attributes; for instance, the ImageNet dataset utilizes WordNet for granularity structure delineation [56]. Besides, pioneers have emerged in the research of automatic granularity discovery and construction, such as community discovery [57]. Researchers can upgrade techniques for the automatic construction of granularity based on the commonalities and specificities proposed in this paper. In addition, considering the distance-based nature of FSDG, the incorporation of optimal transport-based training objectives is another promising direction for further boosting performance [58]. This endeavor would help us gain a deeper understanding of the commonalities and specificities among the samples, as well as the gaps between them and their centroids. Another essential future direction deserving of further exploration is the deeper analysis of the explainability. FSDG seeks to render the black-box features of deep learning more transparent, allowing us to identify certain features responsible for handling commonalities and specificities. Thanks to the numerous breakthroughs in techniques witnessed in recent years, we anticipate a surge of innovation in these promising avenues. These further explorations foster the synergistic optimization of feature structuralization. In summary, our work establishes a strong baseline for exploiting FGDG problems and we believe that the findings presented in this paper warrant further exploration.

### v2

@@ VI Conclusion
In this paper, we assimilate insights from cognitive psychology to advance a feature structuralization framework to tackle the fine-grained domain generalization challenge. FS explicitly embeds structured commonalities and specificities into DL models by integrating data and multi-granularity knowledge. The disentanglement of learned features into common, specific, and confounding segments and their decorrelation, the feature alignment constraints facilitated by three optimization functions, and the prediction calibration term collaboratively contribute to the construction of the FSDG model. Extensive experiments demonstrate a notable enhancement in FGDG performance. Furthermore, the explainability analysis confirms the validity of FS.
Despite FSDG’s systemic efficacy, it brings forth new challenges and prompts intriguing questions. For example, the pre-constructed granularity structure is used to achieve semantic feature alignments. However, some datasets may lack granularity structures, posing challenges to granularity constraints. Notably, one of the prevailing trends in AI research is the pursuit of large-scale datasets, leading to a significant increase in the number of data categories [55]. Such large-scale datasets inherently exhibit hierarchical multi-granularity attributes; for instance, the ImageNet dataset utilizes WordNet for granularity structure delineation [56]. Besides, pioneers have emerged in the research of automatic granularity discovery and construction, such as community discovery [57]. Researchers can upgrade techniques for the automatic construction of granularity based on the commonalities and specificities proposed in this paper. In addition, considering the distance-based nature of FSDG, the incorporation of optimal transport-based training objectives is another promising direction for further boosting performance [58]. This endeavor would help us gain a deeper understanding of the commonalities and specificities among the samples, as well as the gaps between them and their centroids. Another essential future direction deserving of further exploration is the deeper analysis of the explainability. FSDG seeks to render the black-box features of deep learning more transparent, allowing us to identify certain features responsible for handling commonalities and specificities. Thanks to the numerous breakthroughs in techniques witnessed in recent years, we anticipate a surge of innovation in these promising avenues. These further explorations foster the synergistic optimization of feature structuralization. In summary, our work establishes a strong baseline for exploiting FGDG problems, and we believe that the findings presented in this paper warrant further exploration.